OpenStudy (anonymous):

what are the domains x 2 +y 2 + 4x + 6y - 12 = 0

OpenStudy (anonymous):

@dusty

OpenStudy (zpupster):

that is incorrect the input i think should be: x^2 + y^2 + 4x + 6y - 12 = 0 combining and completing the squares- x^2 + 4x + 4) + (y^2 + 6y + 9) = 12 + 9 + 4 * remember to add 4 and 9 to both sides standard circle equation (x+2)^2 + (y+3)^2 -25 = 0 we x+2 is zero when x = -2 and y = -3 will = 0 so our center is (-2,-3) and putting 25 on the right so it is positive and taking the sqre root of 25 = + or - 5 that is our radius so domain ( loosely defined x values) -2 + 5 = 3 and -2 -5 = -7 domain -7 < or = x < or = 3